# Coarse Attribute Prediction with Task Agnostic Distillation for Real World Clothes Changing ReID

### v1

@@ Abstract
This work focuses on Clothes Changing Re-IDentification (CC-ReID) for the real world. Existing works perform well with high-quality (HQ) images, but struggle with low-quality (LQ) where we can have artifacts like pixelation, out-of-focus blur, and motion blur. These artifacts introduce noise to not only external biometric attributes (e.g. pose, body shape, etc.) but also corrupt the model’s internal feature representation. Models usually cluster LQ image features together, making it difficult to distinguish between them, leading to incorrect matches. We propose a novel framework Robustness against Low-Quality (RLQ) to improve CC-ReID model on real-world data. RLQ relies on Coarse Attributes Prediction (CAP) and Task Agnostic Distillation (TAD) operating in alternate steps in a novel training mechanism. CAP enriches the model with external fine-grained attributes via coarse predictions, thereby reducing the effect of noisy inputs. On the other hand, TAD enhances the model’s internal feature representation by bridging the gap between HQ and LQ features, via an external dataset through task-agnostic self-supervision and distillation. RLQ outperforms the existing approaches by 1.6%-2.9% Top-1 on real-world datasets like LaST, and DeepChange, while showing consistent improvement of 5.3%-6% Top-1 on PRCC with competitive performance on LTCC. The code will be made public soon.
@@ 3.1 Proposed Baseline: Clothes Agnostic Base Model (BM)
We adopt one of the most popular open-source CC-ReID model CAL [12] for learning clothes invariant features. CAL focuses on clothing where a clothes classifier predicts clothing labels, while an adversarial loss penalizes correct clothes predictions, thereby removing clothes-relevant information. However, CAL lacks an understanding of pose and gender (Fig. 3). Coupled with its sensitive nature, it is not trivial to integrate additional losses in this architecture. Hence we adopt a two-branch architecture  [14, 4], where Bag-of-tricks (BOT [34]) branch is added to CAL (ResNet-50 backbone split identically after the second block). BOT Branch trains with additional losses and transfers knowledge to the CAL branch via a shared backbone. Concatenation of features from both branches is used during inference $\boldsymbol{f_{ReID}}$ (=$[f_{BOT},f_{CAL}]$). Train-only Identity (ID) classifiers predict ID labels on each branch and KL-divergence on these ID logits ($\mathcal{L}_{ID,KL}$) helps transfer knowledge across the branches indirectly. We use clothes augmented images [61, 19] via body parsing [25] to help the vanilla BOT branch with clothes changing. Collectively these two branches constitute our Base Model which serves as our baseline. Detailed description is provided in Supplmentary. Table 7 indicates that this additional clothes augmentation and two branch structures offer no performance boost, but simply stabilities CAL for additional training modules. Basemodel (CAL and BOT branch are used during inference).
Dataset # of IDs # of Imgs # of Clothes Px. OOF M. Blur Brief Description Evaluation Protocol
@@
LQ Celeb-ReID[15] (TAD-only)$\dagger$ 632 420 20,208 13,978 20,208 - HQ celebrity images -
@@
VLQ DeepChange[49] 450 521 75,083 80,483 1,219 1,336 12 months, 17 views General
@@
Table 1: Dataset, artifact & Evaluation protocol: Px.: denotes pixelation, OOF: denotes out-of-focus blur, M. Blur: denotes motion blur, CC: means clothes changing, SC: means same clothes. $\dagger$: Celeb-reID is chosen as the external dataset for TAD, (not for RLQ evaluation). PRCC and LTCC are real-world LQ datasets, while DeepChange and LaST are Very LQ datasets (VLQ). All artifacts empirically observed.
@@ 3.2 Coarse Attribute Prediction (CAP): External Attributes Features
[…]
@@ 3.3 Task Agnostic Distillation (TAD): Model’s Internal Features
Method Additional Attributes LTCC (%) PRCC (%)
CC General CC SC
[…]
Prev. Work BSGA (BMVC’22) [36] S - - - - 61.8 58.7 99.6 97.3
@@
Figure 4 (bottom) shows TAD. External HQ dataset pre-trains teacher base model and generates synthetic HQ-LQ pairs: $(x_{HQ},x_{LQ})$. We use Celeb-ReID [15] without labels as an external HQ dataset. The frozen teacher only sees $x_{HQ}$ producing normalized CAL and BOT features $[T^{HQ}_{CAL},T^{HQ}_{BOT}]$, while the student sees both $(x_{HQ},x_{LQ})$ generating normalized features as $[S^{HQ}_{CAL},S^{HQ}_{BOT}]$, and $[S^{LQ}_{CAL},S^{LQ}_{BOT}]$. Classifiers produce ID logits $[Y^{\delta}_{CAL,ID},Y^{\delta}_{BOT,ID}]$ and CAL’s clothing logits $[Y^{\delta}_{CAL,CL}]$, where $\delta\in\{HQ,LQ\}$. MSE loss (pink in the figure) distills features, while KL divergence (green in the figure) is used for soft self-supervision
$\mathcal{L}_{D}$ $=\sum_{\delta}^{\{HQ,LQ\}}\|T^{HQ}_{CAL}-S^{\delta}_{CAL}\|_{2}^{2}+\|T^{HQ}_{BOT}-S^{\delta}_{BOT}\|_{2}^{2}$ (2)
$\mathcal{L}_{S}$ $=KL\bigl{(}Y^{HQ}_{CAL,CL}\bigl{\|}Y^{LQ}_{CAL,CL}\bigl{)}+\sum_{\alpha}^{\{CAL,BOT\}}KL\bigl{(}Y^{HQ}_{\alpha,ID}\bigl{\|}Y^{LQ}_{\alpha,ID}\bigl{)}$ (3)
@@ 5 Ablations and Analysis
Following has, BM$+$G: Base Model (BM) w/ only gender classifier. Similarly BM$+$P: w/ only pose branch, BM$+$TAD: w/ only TAD, and BM$+$CAP: RLQ w/o TAD i.e. BM$+$Pose$+$Gender. All ablations are performed with CC protocol for the LTCC dataset (unless mentioned). LTCC is relatively smaller than LaST and DeepChange, making ablation easier. Design choices are highlighted
@@
TAD (Teacher & External Dataset) Table 7 shows variations of TAD for improving robustness against LQ. Base Model with LQ augmentations (BM+LQ Aug.) doesn’t impact the performance. Directly matching student’s HQ-LQ features via MSE loss ($SS_{MSE}$), with teacher $SS_{MSE}$+TAD and without teacher $SS_{MSE}$[NT], offers no performance gain over TAD, while later dropping performance. LTCC+$SS_{MSE}$[NT], $SS_{MSE}$ with noisy HQ-LQ pairs from LTCC instead of clean Celeb-ReID, has low performance. This indicates the need for clean HQ-LQ image pairs, via external dataset. Poor performance without the teacher [NT] indicates a possible feature collapse on the external datasets because of distribution shift, justifying the teacher’s soft supervision for feature separability.
TAD (Task-agnostic) Table 9 shows the use of NTU [30], another HQ dataset, in TAD instead of Celeb-ReID, as a source of synthetic HQ-LQ image pairs. NTU improvement is slightly inferior to Celeb-ReID, possibly because of limited viewpoints. Figure 6 shows random sampling at $\sim$8of Celeb-ReID external dataset maintains consistent performance gain on TAD. This shows that the small amount of data with the teacher’s soft supervision is effective in learning the HQ-LQ relationship, without the need for extensive multi-tasking on the external dataset. These indicate the ReID/task-agnostic nature of the external dataset.
@@ 7 More on CAL Sensitivity problem
Table 11: Contribution of each architectural component on vanilla CAL vs two-branch Base Model (CAP is $+$Gender$+$Pose)
@@
+ Triplet Loss 34.7 (5.4$\downarrow$) 16.6 (1.4$\downarrow$) 71.8 (2.4$\downarrow$) 37.5 (3.3$\downarrow$) 48.6 (6.6$\downarrow$) 49.7 (6.1$\downarrow$) 100 99.8 81.1 (0.6$\downarrow$) 77.0 (2.6$\downarrow$) 81.5 (1.1$\downarrow$) 78.1 (3.2$\downarrow$)

### v2

@@ Abstract
This work focuses on Clothes Changing Re-IDentification (CC-ReID) for the real world. Existing works perform well with high-quality (HQ) images, but struggle with low-quality (LQ) where we can have artifacts like pixelation, out-of-focus blur, and motion blur. These artifacts introduce noise to not only external biometric attributes (e.g. pose, body shape, etc.) but also corrupt the model’s internal feature representation. Models usually cluster LQ image features together, making it difficult to distinguish between them, leading to incorrect matches. We propose a novel framework Robustness against Low-Quality (RLQ) to improve CC-ReID model on real-world data. RLQ relies on Coarse Attributes Prediction (CAP) and Task Agnostic Distillation (TAD) operating in alternate steps in a novel training mechanism. CAP enriches the model with external fine-grained attributes via coarse predictions, thereby reducing the effect of noisy inputs. On the other hand, TAD enhances the model’s internal feature representation by bridging the gap between HQ and LQ features, via an external dataset through task-agnostic self-supervision and distillation. RLQ outperforms the existing approaches by 1.6%-2.9% Top-1 on real-world datasets like LaST, and DeepChange, while showing consistent improvement of 5.3%-6% Top-1 on PRCC with competitive performance on LTCC. Github: https://github.com/ppriyank/RLQ-CGAL-UBD
@@ 3.1 Proposed Baseline: Clothes Agnostic Base Model (BM)
We adopt one of the most popular open-source CC-ReID model CAL [12] for learning clothes invariant features. CAL focuses on clothing where a clothes classifier predicts clothing labels, while an adversarial loss penalizes correct clothes predictions, thereby removing clothes-relevant information. However, CAL lacks an understanding of pose and gender (Fig. 3). Coupled with its sensitive nature, it is not trivial to integrate additional losses in this architecture. Hence we adopt a two-branch architecture  [14, 4], where Bag-of-tricks (BOT [34]) branch is added to CAL (ResNet-50 backbone split identically after the second block). BOT Branch trains with additional losses and transfers knowledge to the CAL branch via a shared backbone. Concatenation of features from both branches is used during inference $\boldsymbol{f_{ReID}}$ ($f_{BOT},f_{CAL}]$). Train-only Identity (ID) classifiers predict ID labels on each branch and KL-divergence on these ID logits ($\mathcal{L}_{ID,KL}$) helps transfer knowledge across the branches indirectly. We use clothes augmented images [61, 19] via body parsing [25] to help the vanilla BOT branch with clothes changing. Collectively these two branches constitute our Base Model which serves as our baseline. Detailed description is provided in Supplmentary. Table 7 indicates that this additional clothes augmentation and two branch structures offer no performance boost, but simply stabilities CAL for additional training modules. Basemodel (CAL and BOT branch are used during inference).
Dataset # of IDs # of Imgs # of Clothes Px. OOF
M.
Blur
Brief
Description
Evaluation
Protocol
@@
LQ
[…]
@@
VLQ
DeepChange[49] 450 521 75,083 80,483 1,219 1,336 12 months, 17 views General
@@
Table 1: Dataset, artifact & Evaluation protocol: Px.: denotes pixelation, OOF: denotes out-of-focus blur, M. Blur: denotes motion blur, $C$: means clothes changing, SC: means same clothes. $\dagger$: Celeb-reID is chosen as the external dataset for TAD, (not for RLQ evaluation). PRCC and LTCC are real-world LQ datasets, while DeepChange and LaST are Very LQ datasets (VLQ). All artifacts empirically observed.
@@ 3.2 Coarse Attribute Prediction (CAP): External Attributes Features
$\mathcal{L}_{Cos}=\big|\text{ Cosine-Similarity}(f_{BOT},f_{POSE})\hskip 1.9919pt\big|$ (1)
@@ 3.3 Task Agnostic Distillation (TAD): Model’s Internal Features
Method
Additional
Attributes
LTCC (%) PRCC (%)
$C$ General $C$ SC
$uparrow R-1\uparrow mAP\uparrow$ $uparrow R-1\uparrow mAP\uparrow$
Prev. Work
BSGA (BMVC’22) [36] S - - - - 61.8 58.7 99.6 97.3
@@
Figure 4 (bottom) shows TAD. External HQ dataset pre-trains teacher base model and generates synthetic HQ-LQ pairs: $(x_{HQ},x_{LQ})$. We use Celeb-ReID [15] without labels as an external HQ dataset. The frozen teacher only sees $x_{HQ}$ producing normalized CAL and BOT features $T^{HQ}_{CAL},T^{HQ}_{BOT}]$, while the student sees both $(x_{HQ},x_{LQ})$ generating normalized features as $S^{HQ}_{CAL},S^{HQ}_{BOT}]$, and $S^{LQ}_{CAL},S^{LQ}_{BOT}]$. Classifiers produce ID logits $Y^{\delta}_{CAL,ID},Y^{\delta}_{BOT,ID}]$ and CAL’s clothing logits $Y^{\delta}_{CAL,CL}]$, where $\delta\in\{HQ,LQ\}$. MSE loss (pink in the figure) distills features, while KL divergence (green in the figure) is used for soft self-supervision
$\mathcal{L}_{D}$ $=\hskip-8.5359pt\sum_{\delta}^{\{HQ,LQ\}}\hskip-5.69046pt\|T^{HQ}_{CAL}-S^{\delta}_{CAL}\|_{2}^{2}+\|T^{HQ}_{BOT}-S^{\delta}_{BOT}\|_{2}^{2}$ (2)
$\mathcal{L}_{S}$ $=\hskip-2.84544ptKL\bigl(Y^{HQ}_{CAL,CL}\bigl\|Y^{LQ}_{CAL,CL}\bigl)+\hskip-17.07182pt\sum_{\alpha}^{\{CAL,BOT\}}\hskip-11.38092ptKL\bigl(Y^{HQ}_{\alpha,ID}\bigl\|Y^{LQ}_{\alpha,ID}\bigl)$ (3)
@@ 5 Ablations and Analysis
Following has, BM++G: Base Model (BM) w/ only gender classifier. Similarly BM++P: w/ only pose branch, BM++TAD: w/ only TAD, and BM++CAP: RLQ w/o TAD i.e. BM++Pose++Gender. All ablations are performed with CC protocol for the LTCC dataset (unless mentioned). LTCC is relatively smaller than LaST and DeepChange, making ablation easier. Design choices are highlighted
@@
[…]
TAD (Task-agnostic) Table 9 shows the use of NTU [30], another HQ dataset, in TAD instead of Celeb-ReID, as a source of synthetic HQ-LQ image pairs. NTU improvement is slightly inferior to Celeb-ReID, possibly because of limited viewpoints. Figure 6 shows random sampling at $\sim$8% of Celeb-ReID external dataset maintains consistent performance gain on TAD. This shows that the small amount of data with the teacher’s soft supervision is effective in learning the HQ-LQ relationship, without the need for extensive multi-tasking on the external dataset. These indicate the ReID/task-agnostic nature of the external dataset.
@@ 7 More on CAL Sensitivity problem
Table 11: Contribution of each architectural component on vanilla CAL vs two-branch Base Model (CAP is ++Gender++Pose)
@@
[…]